# Distilling Reasoning Ability from Large Language Models with Adaptive Thinking

### v5

@@ Abstract
Chain of thought finetuning (cot-finetuning) aims to endow small language models (SLM) with reasoning ability to improve their performance towards specific tasks by allowing them to imitate the reasoning procedure of large language models beyond simply predicting the answers. Most existing cot-finetuning methods adopt a pre-thinking mechanism, allowing the SLM to generate a rationale before answering. In this way, pre-thinking enables SLM to analyze questions but makes answer correctness sensitive to minor errors in rationale. Therefore, we propose a robust post-thinking mechanism to generate answers before rationale. Thanks to this answer-first setting, 1) the answer can escape from the rationale-sensitive problem; 2) the rationale serves as an error amplifier, making SLM focus on learning hard samples; 3) the inferring efficiency can also benefit. Although post-thinking brings many advantages, it may lose the ability to analyze complex questions compared to pre-thinking. Therefore, a plug-and-play adaptive-thinking mechanism is proposed to integrate the merits of pre-thinking and post-thinking, in which a perception module based on soft prompt tuning is introduced to prompt SLM to answer or think first according to the complexity of questions. Extensive experiments are conducted across 12 datasets and 2 language models to demonstrate the effectiveness of the proposed mechanism.
@@ I Introduction
Chain of thought (CoT) [1], as an effective strategy, has been widely used to enhance the capacity of large language models (LLM) [2, 3, 4] on various complex reasoning tasks. However, its usage is still limited since it is observed that such benefits on reasoning capacity brought by CoT only exist in LLM with more than 100B parameters but not in smaller models [5].
@@
To alleviate such problem, prior attempts [6, 7, 8, 9, 10] adopt chain of thought finetuning (cot-finetuning) that is knowledge distillation [11] essentially to transfer reasoning capacity from teacher LLM to student SLM, thereby improving the student SLM’s performance in the vertical field such as complex question answering [12], recommender systems [13], knowledge graphs reasoning [14, 15] and LLM-based Agents [16].
@@
There are two main cot-finetuning paradigms currently. One is the prefixes mechanism [7], and the other is the pre-thinking mechanism [6, 17]. Compared with standard finetuning that only uses the golden answer to train SLM, both of these two cot-finetuning paradigms make SLM imitate the reasoning procedure (also referred to as rationale) of teacher LLM beyond simply predicting the answer during training to elicit the reasoning capacity of student SLM. The difference between them is that the former applies multi-task learning that distinguishes tasks by prefixes, while the latter uses rationale and answer as a sequence generation target. As shown in Fig.1.(b), since rationale and answer are not in the same generation sequence with prefixes mechanism, the contextual association between them is severed, resulting in sub-optimal performance and that rationale mismatches the answer. Thus, pre-thinking is the most developed cot-finetuning mechanism currently.
@@
With the aid of the pre-thinking mechanism, the SLM acquires thinking and reasoning capabilities for complex questions similar to those of LLM. Nevertheless, there is a significant problem (rationale-sensitive problem) with pre-thinking that cannot be overlooked. Specifically, with the next token prediction (NTP) training objective, the rationale-answer generating sequence will lead to the answer generation being overly dependent on the rationale, ultimately making the correctness of answers sensitive to minor errors in rationale. Thus, by looking at the output of SLM to the first question in Fig.1.(c), one can observe that even though the correct information is obtained in the initial part of the rationale, the answer is incorrect due to the subsequent analysis (bold text) deviating from the proper direction. It is worth mentioning that minor errors in rationale are pervasive for SLM.
To solve this problem, we propose a robust post-thinking mechanism. In post-thinking, the rationale-answer generating sequence is adjusted to the answer-rationale sequence. As the answer is generated before rationale, firstly, rationale no longer serves as the key factor in determining the answer but exists as the explanation to the answer. In this way, the student SLM can escape from the adverse effects caused by the minor errors in the rationale, which is intuitively reflected by comparing the outputs of SLM to the first question between Fig.1.(c) and Fig.1.(d). Furthermore, since the answer can be obtained without generating a rationale, post-thinking is more efficient than pre-thinking when inference is conducted. Last but not least, the post-thinking mechanism naturally has an important advantage: the ability to focus on learning hard samples. When SLM yields an incorrect answer, it will seek a weird rationale to justify the erroneous answer. As shown in the rationale of SLM to the second question of Fig.1.(d), SLM tries to justify the incorrect answer from some irrelevant perspectives such as conflicting with family time and being distracted in a public place. This forced interpretation amplifies the model’s errors because the distribution disparity between the rationale output by SLM and the correct rationale will be larger in the incorrect samples compared with that of rationale in samples answered correctly. Higher distribution disparity results in higher training loss for incorrect samples, prompting SLM to focus on learning from hard samples.
@@
Figure 1: Comparison of different cot-finetuning mechanisms. Standard finetuning trains an SLM to generate only the answer to the question. Prefixes train SLM to complete the task of generating answers or rationale under the instruction of different prefixes. Pre-thinking makes SLM think and analyze before answering while Post-thinking lets SLM answer first. Through the proposed perception module, ATM allows SLM to adaptively decide whether to pre-think or post-think based on the complexity of the problem.
@@
However, while post-thinking enables SLM to concentrate on mastering hard samples, it is not feasible to achieve proficiency in tackling highly complex questions merely by escalating the attention of learning rather than changing the way of solving the problem. The second question in Fig1.(d) illustrates a representative hard sample, which differs from the ordinary hard sample in that its higher complexity. For this question, SLM cannot obtain the answer through a single piece of knowledge as it do for the first question in Fig.1.(c). Instead, SLM has to decompose it into two sub-questions: 1) how many Wednesdays in a year and 2) how long it takes to read the Bible 15 times. Obviously, this type of high-complexity hard sample is more suitable for pre-thinking that can break down a question into sub-questions through thinking.
@@
In fact, the mechanism by which humans answer questions is a combination of pre-thinking and post-thinking. People tend to blurt out the answer when a question can be answered directly, while they will think about it when the question is too complicated. Henceforth, motivated by the mechanism human beings employ to answer questions, we introduce the adaptive thinking mechanism (ATM). The ATM evaluates the complexity of the input questions in line with the SLM’s capability, and determines whether to provide an answer directly (post-thinking) or to first think about the question and then output an answer based on the thought (pre-thinking). Specifically, we introduce a perception module within the ATM that gauges the complexity of the input questions and generates a sequence of soft prompt tokens accordingly. These soft prompt tokens are then prepended to the tokens of the input questions to guide SLM on whether the questions necessitate thinking or can be answered directly. Owing to this perception module, the ATM is not only able to preserve the pre-thinking capability to think about and decompose complex questions, but also, aided by post-thinking, can mitigate the rationale-sensitive problem and enhance attention towards learning hard samples. In addition, since ATM is orthogonal to pre-thinking mechanisms, ATM can also be viewed as a plug-and-play enhancement to most methods based on pre-thinking mechanisms.
@@
2) Adaptive thinking mechanism: a plug-and-play mechanism that can perceive the complexity of a question based on SLM’s ability so as to decide whether to answer the question directly. It not only retains the pre-thinking ability to think about and decompose complex questions, but also makes full use of the advantages of the post-thinking mechanism.
@@ II-A Chain of Thought
CoT is first proposed by Wei et al [5] to enhance the model’s reasoning capacity in various complex reasoning tasks. Subsequent attempts optimize CoT in the following three directions.
1) CoT Prompt Construction. There are three methods to construct CoT prompt. The first is the manual CoT prompt that provides different forms of thoughts manually such as natural language [5], programming code [18, 19] or algebras [20]. But manual CoT prompt requires finding task-specific exemplars manually, which limits the scalability and generalization. To reduce the cost of hand-crafted few-shot examples, semi-automatic CoT [21, 22, 23] and automatic CoT [24, 25, 26, 27] are proposed. For instance, Kojima et al[25] adds “Let’s think step by step” after the question so that LLM is able to generate rationale in a zero-shot manner. Auto-CoT [24] generates the demonstrations automatically through clustering and representative exemplars selection.
2) CoT Structural. The chain structure [5, 18] is the most widely used CoT structure. But the scope of rationale exploration is limited by the chain structure. Thus, tree structures [28, 29] and graph structures [30, 31] is proposed. With the help of tree structures and tree search algorithms, LLM gains the ability to explore and backtrack during the reasoning procedure efficiently. Loops and rings in graphs bring about more complex topological relationships compared to trees, which allows LLM to model more complex reasoning.
3) CoT Enhancement Strategy. Since errors existing in CoT often lead to incorrect reasoning steps and answers, there are some attempts proposed to leverage external knowledge [32, 33] or apply ranking and voting mechanisms [34, 35] to enhance the correctness of CoT. External knowledge can supplement knowledge not seen during the pre-training of LLM. And the rank and voting mechanism can reduce the proportion of errors of rationale through ensemble learning.
In our work, we adopt the “Let’s think step by step” to obtain rationale across different tasks due to its convenience.
II-B Knowledge Distillation from LLM
@@
CoT can only improve the reasoning capacity of LLM but not that of small models. Therefore, a lot of works introduce knowledge distillation [11] to transfer the reasoning capacity of LLM to SLM by having SLM study the more informative output of LLM. Different from the previous widely used soft distillation methods that align the distribution of the small model to the large model in the output [36], attention [37], and hidden states [38], the distillation for the reasoning capacity of LLM adopts the hard distillation that aligns the SLM’s generation sequence with that of LLM.
There are two main mechanisms for this hard distillation. 1) prefix mechanism [7]. The prefix mechanism applies task prefixes to distinguish whether the model generates an answer or rationale. Although this mechanism retains the information of the rationale generated from LLM, it separates the contextual connection between the answer and the rationale; 2) pre-thinking mechanism [6, 17, 9]. pre-thinking regards the rationale-answer sequence as the target generation sequence of SLM. However, although multi-step reasoning [39], program-aided distillation [10], counterfactual distillation strategy [40, 41], multiple model integration [42] and the knowledge of negative data [8] are introduced into the pre-thinking mechanism to learn the reasoning capacity of LLM better, the rationale-sensitive problems are still to be resolved.
Thus, we investigate the third mechanism, called post-thinking to distill reasoning capacity from LLM to SLM. Unlike pre-thinking, post-thinking views the answer-rationale sequence as the target generation sequence, which greatly makes the distillation robust to the quality of rationale and SLM focus on learning hard examples. Furthermore, ATM was introduced to fully integrate the advantages of pre-thinking and post-thinking.
II-C Soft Prompt Tuning
@@
For our work, prefix-tuning is applied in the perception module to optimize soft prompt that is used to prompt the SLM to pre-think or post-think. It is worth noting that, unlike the standard soft prompt tuning process where only the parameters related to the soft prompt are optimized, in ATM, the parameters of the language model are continuously optimized as well, because solely optimizing the soft prompt parameters often does not achieve the same effectiveness as finetuning all the parameters for SLM. [48].
@@ III-A Post-Thinking Mechanism
1:  Get $r_{i}$ from LLM using zero-shot prompt “Let’s think step by step
@@
5:  Randomly initialize parameters $\theta_{s}$ and $\theta_{p}$
@@ III-B1 Perception module
the perception module consists of a transformer layer whose hyperparameter configurations are consistent with the transformer layer in SLM.
Firsty, it takes $K$ randomly initialized but learnable token embeddings, termed query tokens $q\in\mathbb{R}^{K\times H}$, and $(d=x^{t}_{i}\oplus a^{t}_{i})\in\mathbb{R}^{L\times H}$ as input, where $H$ and $L$ are the hidden size of SLM and the length of $d$ respectively, $x^{t}_{i}$ and $a^{t}_{i}$ represent the token embedding corresponding to $x_{i}$ and additional feature tokens $a_{i}$ respectively, and $\oplus$ represents the concatenation of the sequence. Various features such as word number and sentence structure of $x_{i}$ contributing to the assessment of the complexity of $x_{i}$ can be included in $a_{i}$. For convenience, only the word number of $x_{i}$ whose specific form is “number of words: num”, is utilized as the feature in our work, where num is the specific word number of the input question. In practical tasks, adaptations to the features in $a_{i}$ could be made to reduce the difficulty of perceiving input questions’ complexity.
Then, the cross-attention module in the transformer layer is implemented to output the soft prompt tokens $p\in\mathbb{R}^{K\times H}$. The specific process can be expressed by the following formula:
@@
where $W_{Q}\in\mathbb{R}^{H\times d_{k}}$, $W_{K}\in\mathbb{R}^{H\times d_{k}}$ and $W_{V}\in\mathbb{R}^{H\times d_{k}}$ refer to the trainable parameter matrices. After the cross-attention module, the useful information in $d$ for perceiving complexity of the question is gathered into $p$ to help the perception module prompt the SLM.
@@ III-B3 Training procedure
TABLE II: The performance on Eecoder-Decoder Language model T5-Large. Accuracy (%) is reported across 12 reasoning tasks with different methods..
@@
The complete training process of the ATM is described in Algorithm 1. Since most of the improvements made by other cot-finetuning methods on pre-thinking can be performed in step 2 and step 8, ATM can further improve the performance of these methods in a plug-and-play manner.

### v6

@@ Abstract
Chain of thought distillation (CoT-distillation) aims to endow small language models (SLM) with reasoning ability to improve their performance towards specific tasks by allowing them to imitate the reasoning procedure of large language models beyond simply predicting the answers. Most existing CoT-distillation methods adopt a pre-thinking mechanism, allowing the SLM to generate a rationale before answering. In this way, pre-thinking enables SLM to analyze questions but makes answer correctness sensitive to minor errors in rationale. Therefore, we propose a robust post-thinking mechanism to generate answers before rationale. Thanks to this answer-first setting, 1) the answer can escape from the rationale-sensitive problem; 2) the rationale serves as an error amplifier, making SLM focus on learning hard samples; and 3) the inferring efficiency can also benefit. Although post-thinking brings many advantages, it may lose the ability to analyze complex questions compared to pre-thinking. Therefore, a plug-and-play adaptive-thinking mechanism is proposed to integrate the merits of pre-thinking and post-thinking, in which a perception module based on soft prompt tuning is introduced to prompt SLM to answer or think first according to the complexity of questions. Extensive experiments are conducted across 12 datasets and 2 language models to demonstrate the effectiveness of the proposed mechanism.
@@ I Introduction
Chain of thought (CoT) [1], as an effective strategy, has been widely used to enhance the performance of large language models (LLM) [2, 3, 4] on various complex reasoning tasks. However, its usage is still limited since it is observed that such benefits on reasoning ability brought by CoT only exist in LLM with more than 100B parameters but not in smaller models [5].
@@
To alleviate such a problem, prior attempts [6, 7, 8, 9, 10] adopt chain of thought distillation (CoT-distillation) to transfer reasoning capacity from teacher LLM to student SLM, thereby improving the student SLM’s performance in the vertical field such as complex question answering [11], recommender systems [12], knowledge graphs reasoning [13, 14] and LLM-based Agents [15].
@@
There are two main CoT-distillation paradigms currently. One is the prefixes mechanism [7], and the other is the pre-thinking mechanism [6, 16]. As shown in Fig.1.(b), since rationale and answer are not in the same generation sequence with prefixes mechanism, the contextual association between them is severed, resulting in sub-optimal performance and that rationale mismatches the answer. Thus, pre-thinking, which generates rationales before answers, becomes the most developed CoT-distillation mechanism.
@@
With the aid of the pre-thinking mechanism, the SLM acquires thinking and reasoning capabilities for complex questions, akin to those exhibited by LLMs. However, a critical issue, termed the rationale-sensitive problem, arises in pre-thinking and cannot be ignored. Specifically, under the next token prediction (NTP) training objective, the rationale-answer generation sequence causes answer generation to become excessively reliant on the rationale, rendering the correctness of answers highly sensitive to minor errors in the rationale. Consequently, as illustrated in the output of the SLM for the first question in Fig.1.(c), although the initial part of the rationale contains correct information, the answer is rendered incorrect due to subsequent analysis (highlighted in bold) deviating from the intended direction. Notably, minor errors in rationale are a pervasive issue for SLM.
To solve this problem, we propose a robust post-thinking mechanism. In the post-thinking mechanism, the rationale-answer generation sequence is restructured into an answer-rationale sequence. Since the answer is generated prior to the rationale, the rationale primarily serves as an explanation for the answer, enabling the student SLM to avoid the adverse effects of minor errors in the rationale. Moreover, as the answer can be generated without the need for a rationale, post-thinking demonstrates higher efficiency compared to pre-thinking during inference. Last but not least, the post-thinking mechanism possesses a notable advantage: its inherent ability to focus on learning from hard samples. When the SLM produces an incorrect answer, it attempts to generate an unusual rationale to justify the error. As illustrated in the rationale provided by the SLM for the second question in Fig.1.(d), the SLM attempts to justify the incorrect answer through irrelevant perspectives, such as conflicts with family time or distractions in a public place. This enforced interpretation in post-thinking, guided by incorrect answers, causes the generated rationale to deviate further from the correct interpretation compared to that produced by pre-thinking, thus leading to greater loss on incorrect samples. This greater loss during training will drive the SLM to concentrate more on learning from the corresponding samples.
@@
Figure 1: Comparison of different CoT-distillation mechanisms. Standard finetuning trains an SLM to generate only the answer to the question. Prefixes train SLM to complete the task of generating answers or rationale under the instruction of different prefixes. Pre-thinking makes SLM think and analyze before answering while post-thinking lets SLM answer first. Through the proposed perception module, ATM allows SLM to adaptively decide whether to pre-think or post-think based on the complexity of the problem.
@@
However, while post-thinking enables the SLM to focus on mastering hard samples, it is impractical to achieve proficiency in addressing highly complex questions solely by intensifying learning attention without altering the problem-solving approach. The second question in Fig1.(d) represents a typical hard sample, distinguished from ordinary hard samples by its higher complexity. For this question, the SLM cannot derive the answer using a single piece of knowledge, as it does for the first question in Fig.1.(c). Instead, the SLM must decompose it into two sub-questions: (1) how many Wednesdays are there in a year, and (2) how long it takes to read the Bible 15 times. Clearly, this type of high-complexity hard sample is better suited for pre-thinking, which can decompose a question into sub-questions through reasoning.
Human cognitive processes can be broadly categorized into fast and slow thinking: the former quickly generates answers based on intuition, while the latter engages in deeper reasoning and analytical processing [17, 18]. When answering questions, humans select between fast and slow thinking depending on factors such as prior experience, environmental context, and question complexity, aiming to optimize both efficiency and effectiveness. Similarly, post-thinking and pre-thinking mirror characteristics analogous to fast and slow thinking, respectively. Building on this cognitive paradigm, we propose the adaptive thinking mechanism (ATM), which dynamically selects between pre-thinking and post-thinking based on the SLMs’ capabilities and the complexity of a given question.
@@
Specifically, we incorporate a perception module into the ATM, which evaluates the complexity of input questions and generates a sequence of soft prompt tokens accordingly. These soft prompt tokens are then prepended to the input tokens to guide the SLM in determining whether the questions require reasoning or can be answered directly. Owing to this perception module, the ATM not only retains the pre-thinking capability to reason through and decompose complex questions but also, aided by post-thinking, mitigates the rationale-sensitive problem and enhances the focus on learning hard samples. Furthermore, as the ATM is orthogonal to pre-thinking mechanisms, it can also serve as a plug-and-play enhancement for most methods based on pre-thinking.
@@
2) Adaptive thinking mechanism: a plug-and-play framework that adaptively selects between pre-thinking and post-thinking based on the complexity of the question, effectively integrating the strengths of both approaches.
@@ II-A Chain of Thought Distillation
CoT is first proposed by Wei et al [5] to enhance the model’s reasoning capacity in various complex reasoning tasks. Subsequent attempts optimize CoT in the following three directions: constructing more effective CoT prompt [19, 20, 21, 22, 23], modeling more complex reasoning with tree [24, 25] or graph [26, 27] structure, enhancing reasoing with external knowledge, tools, and post-processing strategy [28, 29, 30, 31].
@@
However, CoT can only improve the reasoning capacity of LLM but not that of SLM [5]. Therefore, a lot of works introduce knowledge distillation [32] to transfer the reasoning capacity of LLM to SLM by having SLM study the more informative output of LLM. Different from the previous widely used soft distillation methods that align the distribution of the small model to the large model in the output [33], attention [34], and hidden states [35], the distillation for the reasoning capacity of LLM adopts the hard distillation that aligns the SLM’s generation sequence with that of LLM.
There are two main mechanisms for this hard distillation.
II-A1 Prefix mechanism
Distilling step by step [7] is a pioneer of the prefix mechanism, which applies task prefixes to distinguish whether the model generates an answer or rationale. Based on it, Mind’s Minor [36] introduces another self-evaluation task to further improve model performance and Mint [37] explores the effect of prefix mechanism on mathematical reasoning. Although these methods retain the information of the rationale generated from LLM, it separates the contextual connection between the answer and the rationale, resulting in a mismatch between them. To alleviate the mismatch, MMI [38] maximizes the mutual information between the answer and the rationale. However, since the problem of not being in the same context has not been fundamentally solved, the mismatch problem is still quite significant.
II-A2 Pre-thinking mechanism
Basic pre-thinking treats the rationale-answer sequence as the target generation sequence for SLM [6, 16]. Numerous methods have been proposed to enhance the performance of the basic pre-thinking mechanism. Multi-step reasoning [39], MD [40], and UniCoTT [41] combine various forms of data (e.g., different CoT structures) for training. DISCO [42], SCOTT [43], and MVC-CD [44] incorporate counterfactual distillation strategies into pre-thinking. Pad [10] and MD [40] integrate programming-aid capabilities. SGFT [45] explores the effect of distilling solution guidance instead of complete rationales. KPOD [46] proposes a progressive distillation strategy to help SLMs better comprehend the rationale provided by LLMs. The incorporation of negative data [8] and multiple model integration [47] has also been shown to enhance the pre-thinking mechanism.
However, while the aforementioned methods improve the ability of SLMs to learn the reasoning capacity of LLMs, the rationale-sensitive problems remain unresolved. Thus, we investigate the third mechanism, called post-thinking to distill reasoning capacity from LLM to SLM. Furthermore, ATM is introduced to fully integrate the advantages of pre-thinking and post-thinking. It is worth noting that there are several human-inspired dual-system thinking frameworks [48, 49, 50, 51, 52], comprising fast and slow thinking, to enhance the reasoning capabilities of language models. Compared to ATM, these methods exhibit two limitations: 1) their fast thinking component still relies on pre-generated rationales, albeit shorter in length; and 2) the SLM mimics external rules or LLMs’ predictions, rather than being informed by its internal capabilities to chosse thinking mode, resulting in suboptimal performance.
II-B Soft Prompt Tuning
@@
For our work, prefix-tuning is applied in the perception module to optimize the soft prompt that is used to prompt the SLM to pre-think or post-think. It is worth noting that, unlike the standard soft prompt tuning process where only the parameters related to the soft prompt are optimized, in ATM, the parameters of the language model are continuously optimized as well, because solely optimizing the soft prompt parameters often does not achieve the same effectiveness as finetuning all the parameters for SLM [58].
@@ III-A Post-Thinking Mechanism
1: Get $r_{i}$ from LLM using zero-shot prompt “Let’s think step by step why the answer is $y_{i}$
@@
5: Initialize parameters $\theta_{s}$ and $\theta_{p}$
@@ III-B1 Perception module
The perception module consists of a transformer attention layer whose hyperparameter configurations are consistent with the transformer layer in SLM.
Firsty, it takes $K$ randomly initialized but learnable token embeddings, termed query tokens $q\in\mathbb{R}^{K\times H}$, and $(d=x^{t}_{i}\oplus a^{t}_{i})\in\mathbb{R}^{L\times H}$ as input, where $H$ and $L$ are the hidden size of SLM and the length of $d$ respectively, $x^{t}_{i}$ and $a^{t}_{i}$ represent the token embedding corresponding to $x_{i}$ and additional feature tokens $a_{i}$ respectively, and $\oplus$ represents the concatenation of the sequence. $x_{i}$ inherently contains implicit complexity-related features, such as word count, linguistic complexity, readability, and the number of reasoning steps required. $a_{i}$ is intended to reduce the learning difficulty of the perception module during training by providing additional explicit information. For convenience, only the word number of $x_{i}$ whose specific form is “number of words: num”, is utilized as the feature in our main experiments, where num is the specific word number of the input question. In practical tasks, adaptations to the features in $a_{i}$ could be made to reduce the difficulty of perceiving input questions’ complexity.
Then, the cross-attention module is implemented to output the soft prompt tokens $p\in\mathbb{R}^{K\times H}$. The specific process can be expressed by the following formula:
@@
where $W_{Q}\in\mathbb{R}^{H\times H}$, $W_{K}\in\mathbb{R}^{H\times H}$ and $W_{V}\in\mathbb{R}^{H\times H}$ refer to the trainable parameter matrices. After the cross-attention module, the useful information in $d$ for perceiving complexity of the question is gathered into $p$ to help the perception module prompt the SLM.
@@ III-B3 Training procedure
TABLE II: The performance on Encoder-Decoder Language model T5-Large. Accuracy (%) is reported across 12 reasoning tasks with different methods.
@@
The complete training process of the ATM is described in Algorithm 1. Since most of the improvements made by other CoT-distillation methods on pre-thinking can be performed in step 2 and step 8, ATM can further improve the performance of these methods in a plug-and-play manner.